# More Agents Is All You Need

### v1

@@ 5.1 Generalizability
Table 2 and Figure 3 show that our method generally enhances performance across all tasks and LLMs by increasing the ensemble size. Specifically, in arithmetic reasoning tasks, the accuracy gains range from $12\%$ to $24\%$ on the GSM8K and from $6\%$ to $10\%$ on the MATH. In general reasoning tasks, the accuracy gains range from $1\%$ to $4\%$ on the Chess and from $5\%$ to $11\%$ on the MMLU. In code generation task, the accuracy gains range from $4\%$ to $9\%$ on HumanEval. Surprisingly, our method enables a smaller LLM to outperform a larger counterpart by simply scaling up the ensemble size. For instance, the enhanced Llama2-13B model achieves $59\%$ accuracy on the GSM8K dataset, outperforming the Llama2-70B model, which scores $54\%$.
@@ 5.2 Compatibility
Table 3 shows that by integrating our method with other methods, the performance can be further improved across different LLMs and tasks, despite these methods have different implementations. To be specific, in arithmetic reasoning tasks, our method enhances these methods to further improvement, yielding increases between $10\%$ and $21\%$ on the GSM8K dataset, and between $1\%$ and $15\%$ on the MATH dataset. In general reasoning tasks, integration with other methods generally achieves performance gains ranging from $1\%$ to $13\%$ in the Chess task and from $1\%$ to $11\%$ in the MMLU task. In code generation task, when combined with other methods, gains range from $2\%$ to $7\%$. However, two notable exceptions are observed when integrated with the debate method with the Llama2-13B and Llama2-70B models, which result in failed cases. This failure in performance is attributed primarily to the noise generated by referencing the answers of other agents during the debate process. The synthesized responses, which incorporate input from multiple agents, disrupt the coherence of the code logic, leading to the observed performance degradation. All accuracy curves are provided in the Appendix B.
@@ 5.4 Robustness
We conduct ablation studies to evaluate the impact of changes in various hyperparameters on the final performance. The experiment is conducted by altering the temperature T (Ficler & Goldberg, 2017) and the nucleus probability p (Radford et al., 2019), using the GPT-3.5-Turbo model over an average of 20 runs. As shown in Figure 4, scaling up ensemble size improves the LLM performance consistently across different tasks, despite the variation of these hyperparameters.
@@ 6.2 Inherent Difficulty
Figure 6: (Left) The relative performance gains increase and then decrease with rising inherent difficulty. (Middle) The relative performance gains increase with the number of steps. (Right) The absolute performance increases with the prior probability. We analyze each dimension by fixing the other two dimensions.
@@ 6.3 Number of Steps
Property 2.2: Sampling-and-voting increases the performance for each step. We conduct a fine-grained analysis for each step of a given task. We explicitly prompt the language model to output the result of each step. Subsequently, we utilize sampling-and-voting at each step to derive the answer for that step. Figure 7 (left) shows that although each step has equal inherent difficulty, the accumulation of errors from previous steps lead to a decrease in accuracy as the number of steps increases. However, our method mitigates the performance decrease encountered with increasing steps.
Figure 7: (Left) Our method increases the performance for each step. Blue bars show the accuracy of various steps for a single sample, and orange bars show the gains for 40 samples. (Middle) Step-wise sampling-and-voting can further enhance the performance across different levels of inherent difficulty. (Right) Hierarchical sampling-and-voting can further enhance the performance with homogeneous and heterogeneous model combinations.
Derivation. Based on Property 2, We propose step-wise sampling-and-voting can further enhance the performance.
Step-wise sampling-and-voting initially prompts the LLM to decompose the task into multiple steps. It then proceeds with multi-round iterations to produce the final result. In each round, the process begins by selecting a current unprocessed step and using sampling-and-voting to determine the result of that step. Subsequently, it uses the result to update the task. This iterative process is repeated multiple times until the last step is processed. To evaluate the performance of step-wise sampling-and-voting, we fix $S=8$ and $K=4$, and tune $I$ from $100$ to $400$. Figure 7 (middle) shows that compared to simple sampling-and-voting, step-wise sampling-and-voting yields greater improvements. e.g., we see $15\%$-$42\%$ gains, which increase with inherent difficulty.
@@ 6.4 Prior Probability
Derivation. Based on Property 3, we propose hierarchical sampling-and-voting can further enhance the performance.
@@ 7 Conclusions and Future Work
The results indicate that our simple sampling-and-voting method for instantiating agents can generally improve the performance of LLMs by increasing the ensemble size. Importantly, this method is orthogonal to different existing methods, which can lead to further improvements when combined with them.
@@
Considering that each input remains the same when we increase the number of agents, the sampling phase can be optimized to reduce the cost. Nonetheless, such a challenge of escalating costs commonly exist in works requiring multiple LLM calls (Wang et al., 2023b; Du et al., 2023). We leave it as a future work to optimize.
Impact Statement
This paper introduces a simple method designed to enhance the performance of Large Language Models (LLMs). While the proposed method aims to improve the efficacy of LLMs in various tasks, it is necessary to acknowledge the potential risks. LLMs can sometimes produce outputs that, while plausible, may be factually incorrect or nonsensical. Such hallucinations can lead to the misguidance of decision-making processes and the propagation of biases. These concerns are particularly acute in the context of critical decision-making scenarios, where the accuracy and reliability of information are paramount. The broader adoption of LLMs, without adequate safeguards against these risks, could exacerbate these issues. Therefore, it is crucial to continue developing mechanisms to mitigate the potential adverse effects of LLM hallucinations to ensure that the deployment of these powerful models is both responsible and beneficial.
@@ A.2 Experiments on Arithmetic Reasoning Tasks
For the implementation of the sampling-and-voting method to arithmetic reasoning tasks within the GSM8K and MATH datasets, we execute the initial sampling phase by Algorithm 15. Samples are extracted from the responses by matching “boxed{{X}}”, where “X” denotes a numerical value or a mathematical expression.
@@ A.3 Experiments on General Reasoning Tasks
For general reasoning tasks, as encountered in the MMLU and Chess datasets, the sampling-and-voting method is applied following Algorithm 15 during the sampling phase. Samples are extracted by matching the pattern “(X” or “(X)”, where “X” corresponds to the options A, B, C, or D in MMLU task and the chessboard position in Chess task.
@@ A.4 Experiments on Code Generation Task
In the code generation task, we apply the sampling-and-voting method to produce Python code using the HumanEval dataset. During the sampling phase, we extract Python code snippets from the model’s responses.
In the voting phase, we compute the BLEU score using sacreBLEU (Post, 2018) to evaluate the similarity between each of the generated samples. The sample with the highest cumulative BLEU score is selected as the final answer. This method ensures that the final output is the most representative and accurate piece of code as determined by consensus through similarity scoring among the samples.

### v2

@@ 5.1 Generalizability
Table 2 and Figure 3 show that our method generally enhances performance across all tasks and LLMs by increasing the ensemble size. Specifically, in arithmetic reasoning tasks, the accuracy gains range from $12\%$ to $24\%$ on the GSM8K and from $6\%$ to $10\%$ on the MATH. In general reasoning tasks, the accuracy gains range from $1\%$ to $4\%$ on the Chess and from $5\%$ to $11\%$ on the MMLU. In code generation task, the accuracy gains range from $4\%$ to $9\%$ on HumanEval. Surprisingly, our method enables a smaller LLM to outperform a larger counterpart by simply scaling up the ensemble size. For instance, the enhanced Llama2-13B model achieves $59\%$ accuracy on the GSM8K dataset, outperforming the Llama2-70B model, which scores $54\%$. Additional statistical results are presented in Appendix B.2.
@@ 5.2 Compatibility
Table 3 shows that by integrating our method with other methods, the performance can be further improved across different LLMs and tasks, despite these methods have different implementations. To be specific, in arithmetic reasoning tasks, our method enhances these methods to further improvement, yielding increases between $10\%$ and $21\%$ on the GSM8K dataset, and between $1\%$ and $15\%$ on the MATH dataset. In general reasoning tasks, integration with other methods generally achieves performance gains ranging from $1\%$ to $13\%$ in the Chess task and from $1\%$ to $11\%$ in the MMLU task. In code generation task, when combined with other methods, gains range from $2\%$ to $7\%$. However, two notable exceptions are observed when integrated with the debate method with the Llama2-13B and Llama2-70B models, which result in failed cases. This failure in performance is attributed primarily to the noise generated by referencing the answers of other agents during the debate process. The synthesized responses, which incorporate input from multiple agents, disrupt the coherence of the code logic, leading to the observed performance degradation. All accuracy curves are provided in the Appendix B.1.
@@ 5.3 Effectiveness
Method +Ours GPT-3.5 70B 13B Overall COT Wei et al. (2022) 2.8 3.6 3.6 3.3 ZS-COT Kojima et al. (2022) 2.8 2.4 3 2.7 SPP Wang et al. (2023c) 4.6 3.6 3.8 4 Debate Du et al. (2023) 3.8 4.4 5 4.4 Reflection Shinn et al. (2023) 3 4.0 3 3.3 Ours 2.6 2.6 2.2 2.5
@@ 5.4 Robustness
We conduct ablation studies to evaluate the impact of changes in various hyperparameters on the final performance. The experiment is conducted by altering the temperature T Ficler & Goldberg (2017) and the nucleus probability p Radford et al. (2019), using the GPT-3.5-Turbo model over an average of 20 runs. As shown in Figure 4, scaling up ensemble size improves the LLM performance consistently across different tasks, despite the variation of these hyperparameters.
5.5 Token usage
We record the token usage for different methods, with the Table 5 presenting the token usage for a single agent. Given that our method scales up by increasing the ensemble size, the token usage increases proportionally with the number of agents or when combined with other methods. When addressing specific tasks, one can trade a higher token budget for improved performance. More details are presented in Appendix B.3
Table 5: Token usage of GPT-3.5
Method GSM8K MATH Chess MMLU HumanEval Ours (single agent) 235 ± 54 326 ± 131 138 ± 14 247 ± 91 495 ± 120 COT Wei et al. (2022) 261 ± 63 360 ± 134 284 ± 76 330 ± 110 330 ± 116 ZS-COT Kojima et al. (2022) 228 ± 56 305 ± 122 132 ± 12 230 ± 92 305 ± 132 SPP Wang et al. (2023c) 341 ± 85 471 ± 161 363 ± 35 428 ± 123 501 ± 125 Reflection Shinn et al. (2023) 214 ± 45 281 ± 96 146 ± 13 218 ± 80 338 ± 110
@@ 6.2 Inherent Difficulty
Figure 6: (Left) The relative performance gains increase and then decrease with rising inherent difficulty. (Middle) The relative performance gains increase with the number of steps. (Right) The absolute performance increases with the prior probability. We analyze each dimension by fixing the other two dimensions. The error bars represent the standard error.
@@ 6.3 Number of Steps
Property 2.2: Agent Forest increases the performance for each step. We conduct a fine-grained analysis for each step of a given task. We explicitly prompt the language model to output the result of each step. Subsequently, we utilize Agent Forest at each step to derive the answer for that step. Figure 7 (left) shows that although each step has equal inherent difficulty, the accumulation of errors from previous steps lead to a decrease in accuracy as the number of steps increases. However, our method mitigates the performance decrease encountered with increasing steps.
Figure 7: (Left) Our method increases the performance for each step. Blue bars show the accuracy of various steps for a single sample, and orange bars show the gains for 40 samples. (Middle) Step-wise Agent Forest can further enhance the performance across different levels of inherent difficulty. (Right) Hierarchical Agent Forest can further enhance the performance with homogeneous and heterogeneous model combinations.
Derivation. Based on Property 2, we propose Step-wise Agent Forest, which can further enhance the performance.
Step-wise Agent Forest initially prompts the LLM to decompose the task into multiple steps. It then proceeds with multi-round iterations to produce the final result. In each round, the process begins by selecting a current unprocessed step and using Agent Forest to determine the result of that step. Subsequently, it uses the result to update the task. This iterative process is repeated multiple times until the last step is processed. To evaluate the performance of Step-wise Agent Forest, we fix $S=8$ and $K=4$, and tune $I$ from $100$ to $400$. Figure 7 (middle) shows that compared to Agent Forest, Step-wise Agent Forest yields greater improvements. e.g., we see $15\%$-$42\%$ gains, which increase with inherent difficulty.
@@ 6.4 Prior Probability
Derivation. Based on Property 3, we propose Hierarchical Agent Forest can further enhance the performance.
@@ 7 Conclusions and Future Work
The results indicate that Agent Forest can generally improve the performance of LLMs by increasing the ensemble size. Importantly, this method is orthogonal to different existing methods, which can lead to further improvements when combined with them.
@@
Considering that each input remains the same when we increase the number of agents, the sampling phase can be optimized to reduce the cost. Nonetheless, such a challenge of escalating costs commonly exists in works requiring multiple LLM calls Wang et al. (2023b); Du et al. (2023). This aligns with recent findings as discussed in Kapoor et al. (2024) that current benchmarks focus narrowly on accuracy, leading to costly agents. Additionally, the lack of adequate holdout sets and standardization in evaluation practices further complicates the development of cost-effective agents. Addressing these issues can enhance the practical efficiency of AI agents in real-world scenarios. We leave it as a future work to optimize.
@@ A.2 Experiments on Arithmetic Reasoning Tasks
For the implementation of Agent Forest to arithmetic reasoning tasks within the GSM8K and MATH datasets, we execute the initial sampling phase by Algorithm 15. Samples are extracted from the responses by matching “boxed{{X}}”, where “X” denotes a numerical value or a mathematical expression.
@@ A.3 Experiments on General Reasoning Tasks
For general reasoning tasks, as encountered in the MMLU and Chess datasets, Agent Forest is applied following Algorithm 15 during the sampling phase. Samples are extracted by matching the pattern “(X” or “(X)”, where “X” corresponds to the options A, B, C, or D in MMLU task and the chessboard position in Chess task.
@@ A.4 Experiments on Code Generation Task
In the code generation task, we apply the Agent Forest method to produce Python code using the HumanEval dataset. During the sampling phase, we extract Python code snippets from the model’s responses.
In the voting phase, we compute the BLEU score using sacreBLEU Post (2018) to evaluate the similarity between each of the generated samples. The sample with the highest cumulative BLEU score is selected as the final answer. This method ensures that the final output is the most representative and accurate piece of code as determined by consensus through similarity scoring among the samples.
@@ B.1 Additional Accuracy Curves
B.2 Statistical Test Results
In this section, we performed a one-way ANOVA to determine if there are significant differences in accuracy across different ensemble sizes (1, 10, 20, 30, 40), across 10 independent runs. The detailed p-values are provided in Table 7, where all p-values are less than 0.05, demonstrating that the differences in accuracy between ensemble sizes are significant.
Table 7: Statistical test results (p-values) of different LLMs on different datasets. One-way ANOVA is performed where ensemble sizes are (1, 10, 20, 30, 40).
LLMs GSM8K MATH Chess MMLU HumanEval Llama2-13B 1.56e-44 9.11e-32 1.97e-18 7.86e-38 5.46e-8 Llama2-70B 1.93e-43 2.16e-29 1.09e-9 1.06e-26 5.24e-7 GPT-3.5-Turbo 6.03e-30 9.66e-39 4.62e-24 2.46e-39 7.47e-13
Figure 14: Token usage vs. accuracy.
B.3 Token Usage vs. Accuracy
Figure 14 shows the tradeoff between token budget and accuracy, where the x-axis represents the token budget and the y-axis represents the accuracy.